OpenStudy (anonymous):

Find cosAcosB + sinAsinB + tanAtanB. If A = 3^(1/2) + 3^(1/2) + 3^(1/2)....till infinity and B = 2*pi*r*3^(1/2)

OpenStudy (anonymous):

\[\cos(A) \cos(B) + \sin(A)\sin(B) = \cos(A-B)\]

OpenStudy (anonymous):

My brain has got rusty.. Unable to think beyond this. :(

OpenStudy (anonymous):

@mukushla .. Benim yoldashim ..